# Improving Human-AI Coordination through Online Adversarial Training and Generative Models

### v3

@@ Abstract
Being able to cooperate with new people is an important component of many economically valuable AI tasks, from household robotics to autonomous driving. However, generalizing to novel humans requires training on data that captures the diversity of human behaviors. Adversarial training is a promising method that allows dynamic data generation and ensures that agents are robust. It creates a feedback loop where the agent’s performance influences the generation of new adversarial data, which can be used immediately to train the agent. However, adversarial training is difficult to apply in a cooperative task; how can we train an adversarial cooperator? We propose a novel strategy that combines a pre-trained generative model to simulate valid cooperative agent policies with adversarial training to maximize regret. We call our method GOAT: Generative Online Adversarial Training. In this framework, the GOAT dynamically searches the latent space of the generative model for coordination strategies where the learning policy—the Cooperator agent—underperforms. GOAT enables better generalization by exposing the Cooperator to various challenging interaction scenarios. We maintain realistic coordination strategies by keeping the generative model frozen, thus avoiding adversarial exploitation. We evaluate GOAT with real human partners, and the results demonstrate state-of-the-art performance on the Overcooked benchmark, highlighting its effectiveness in generalizing to diverse human behaviors.  [^1] Please check out our live interactive demo, game play videos, and training code at our website. We also provide a runnable Google Colab for the cooperative matrix game.
@@ 1 Introduction
Figure 1: Adversarial training framework for cooperative agents. There are two main components: (Left) A generative model encodes simulated agents into a latent space to learn diverse agent strategies, which are then used to simulate different types of agents. (Right) GOAT samples new agents to maximize the Cooperator agent’s regret, defined as the performance gap between self-play (GOAT sampled agents playing with themselves) and cross-play (the same GOAT sampled agent playing with the Cooperator). The key idea is that because the adversarial objective does not directly update the weights of the generative model, it can only simulate valid, cooperative partners that do not engage in self-sabotage. But by applying adversarial training to search over the policies that can be generated by the model, we can expose the Cooperator to a curriculum of challenging training partners, ensuring it is robust to playing with diverse human partners at test time.
@@
In multi-agent human-AI cooperation [3], training a cooperative agent to generalize across diverse human behaviors remains a formidable challenge. Distribution shift, limited training data, and complex human decision-making make training these systems difficult. Without adequate coverage of all possible human interactions, AI agents learn brittle policies and struggle to generalize to novel, unseen behaviors in real human cooperation partners.
@@
Adversarial training could provide a computationally efficient solution for training robust Cooperators by exposing both the Cooperator agent’s limitations and underexplored areas of the partner strategy space. It can be used to automatically generate a curriculum of challenging tasks targeted to the learner’s weaknesses, resulting in a more robust policy [6]. Yet despite its potential, adversarial training within cooperative AI introduces unique challenges; what does it mean to train an adversarial cooperator? In a single-agent task, the Adversary has the sole objective to challenge the learning agent as much as possible by minimizing its performance. However, in a collaborative task, the adversarial agent has to find weaknesses in the learner while still simulating a realistic cooperative partner. Suppose the Adversary is successful in finding weaknesses. In that case, it can become overly combative, resorting to strategies that sabotage the cooperation task, and is thus no longer a valid simulation of what a human cooperation partner might do. The challenge lies in simulating realistic yet challenging partners that maintain cooperative intentions while simultaneously correcting weaknesses in the AI agent’s policy.
For single-agent tasks, techniques like PAIRED [6] constrain the Adversary so that it can only generate valid tasks for which there is a solution. This is accomplished by maximizing the learning agent’s regret, which is the performance gap between the optimal agent and the learning agent. The regret objective proves effective as the learning agent is actively challenged with a curriculum of increasingly difficult tasks, however, each task remains feasible.
In the cooperative setting, attempts have been made to create a regret-like objective for training more diverse simulated partner populations. Charakorn et al. [4] proposes a training objective to make a population of mutually incompatible partners, based on balancing the difference between two terms, self-play (SP) and cross-play (XP) performance. SP is the score a policy achieves when paired with itself, and XP is the score a policy receives when cooperating with a different policy. Adding the SP term ensures each policy can cooperate effectively (by achieving high scores with itself), while minimizing XP encourages the emergence of distinct strategies between policies. However, naively minimizing XP can once again lead to policies learning to sabotage the game. They can learn a handshaking protocol, such that if their partner fails to complete the handshake correctly, they become adversarial and try to strategically sabotage the game to minimize the XP score. While attempts have been made to add additional objectives to counterbalance this effect Sarkar et al. [28], they have not been able to address the issue that minimizing cross-play inherently incentivizes uncooperative behavior.
@@
How can we ensure that the adversarial agent can only generate valid cooperative behaviors that do not engage in self-sabotage but can still fully explore the space of challenging partner strategies? We propose a novel adversarial training method to achieve these criteria. We first pre-train a generative model of partner strategies on valid, diverse cooperation data, following Liang et al. [18]. We then place this generative model into an adversarial training loop, where we train an Adversary to search for embedding vectors that, when passed through the generative model, decode into simulated partners that maximize the Cooperator’s regret (as approximated via SP-XP). Because we only update the weights of the Adversary and not the generative model, we ensure that all generated partners are valid and cooperative, thereby addressing the intentional sabotage seen in adversarial training. In this algorithm, the Adversary agent is motivated to maximize the regret, searching for valid coordination strategies where the Cooperator agent underperforms. In contrast, the Cooperator attempts to adapt to the Adversary and minimize the regret. Unlike previous adversarial approaches that risk learning degenerate strategies, this cyclic training method provides a natural curriculum for the Cooperator agent to train robustly and generalize to diverse partner strategies.
@@
We evaluate GOAT on a toy environment, One-Step Cooperative Matrix Game (CMG), and Overcooked [3], a popular human-AI zero shot coordination benchmark, which is a cooperative cooking game where agents must collaborate to prepare soup recipes successfully. On the CMG toy environment, we find that GOAT discovers and exploits all the regions of reward space with multiple maxima and minima, while other baselines [43, 20, 28] get stuck or improperly assign weights to certain modes. We tested GOAT live against real, novel human partners and found that it significantly improves cooperation performance. GOAT achieves improved performance when evaluated against 5 competitive baselines [43, 28, 18, 32, 3].
@@
  • We introduce a novel adversarial training method for zero-shot coordination, which uses an Adversary to search over the space of partner policies that a generative model can simulate. Because the generative model can only simulate valid cooperative partners, it ensures the adversarial training process generates challenging but realistic cooperative partners.
  • We train both the Adversary and Cooperator agents online in a common payoff game. This results in a dynamic curriculum where the Adversary actively adopts strategies tailored to challenge the Cooperator policy’s current weaknesses, while the Cooperator learns to effectively correct them.
  • We perform a real-time evaluation with novel human partners, using the popular Overcooked benchmark [3]. As compared to recent high-scoring techniques, we show that our method leads to the best cooperation performance with diverse people.
@@ 2 Related Work
Zero-Shot coordination (ZSC). Human-AI coordination requires training cooperative agents that can generalize to interactions with different humans in zero-shot settings [13, 17]. Training effective Cooperator agents requires navigating a vast coordination behavior space. Behavior cloning [3] is limited by expensive and often scarce human data. Self-Play (SP) [27, 30, 31, 21] is an automatic training algorithm that does not need human data but converges to rigid policies that generalize poorly to novel human partners. Population-Based Training (PBT) [14, 35, 22, 15] has become a popular automatic training paradigm for ZSC, in which a Cooperator policy is trained using a large population of simulated partners [12, 32, 20, 4, 43, 28].
@@
Population diversity objectives. In an effort to ensure the simulated agent population covers diverse human strategies, various methods [12, 22, 39, 20, 43, 4, 28, 24] improve PBT by optimizing the diversity of the agents in the population. The population can be created by reward-shaping [33, 42], manual design [11, 37, 40], and quality diversity [2, 23, 39, 9]. However, these methods typically use domain knowledge and are not general and scalable. Statistical methods [12, 8, 22, 7, 20, 43] automate the diverse population generation but Sarkar et al. [28] demonstrated that trajectory distribution methods can result in minor variations in the policies rather than behaviorally different policies.
@@
Another line of work maximizes diversity by training self-play policies while minimizing cross-play rewards with the policies in the population [5, 4, 28]. However, agents might learn to sabotage the cross-play game while retaining high self-play performance with handshaking protocols. To prevent overly adversarial behavior, Sarkar et al. [28] introduces mixed-play, balancing self-play and cross-play by randomly sampling initial game states from both. However, such a solution still leads to rigid policies that do not generalize. In this work, we use online adversarial training to adapt to and exploit the cooperators’ weakness, rather than being limited by the coordination strategies in a pre-built population. And since we only sample from a frozen, pre-trained generative model, we prevent the adversary from learning handshaking protocols or sabotaging behavior.
@@
Generative models for Cooperation. Derek and Isola [7], Wang et al. [36], Liang et al. [18] show that generative models can be trained on policies to simulate diverse and valid collaborative policies. Recently, Liang et al. [18] proposed training a Variational Autoencoder [16], on simulated cooperative agent trajectories to generate synthetic training data to improve human-AI coordination. This approach outperforms previous methods because the generative model enables sampling novel partner strategies that interpolate between or combine strategies from the data. GOAT leverages this approach in an adversarial training loop.
Adversarial training. Adversarial training [34, 5, 10, 26] has emerged as one of the most effective methods for improving generalization to novel scenarios. Methods like PAIRED [6] generate an automatic curriculum where an Adversary creates challenging yet feasible scenarios for the learning agent. This method is effective because it is based on a regret objective that incentivizes the Adversary agent to produce increasingly challenging but feasible tasks, addressing a crucial point: constraining the Adversary from minimizing performance to the point where the learner cannot receive any reward and training becomes stalled. We propose using regret-based adversarial training to focus on presenting partner strategies that target weaknesses in the Cooperator agent, but are still valid cooperative strategies. Our results demonstrate that the use of a regret-based objective is critical to the success of our method.
@@ 3 Preliminaries
Two-player Markov game: We consider the two-player Markov game [19] as a tuple $\bigl{(}\,\mathcal{S},\,\mathcal{A}_{1},\,\mathcal{A}_{2},\,r,\;\mathcal{T},\;\gamma\bigr{)}$ where $\mathcal{S}$ is the state space shared by all the agents. $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ denote the action space of agents such that at each time step in the environment, each agent selects one action $a_{1,t}\in\mathcal{A}_{1}$, $a_{2,t}\in\mathcal{A}_{2}$ sampled from their own action space according to their policies $\pi_{1}:\mathcal{S}\to\Delta(\mathcal{A}_{1})$ and $\pi_{2}:\mathcal{S}\to\Delta(\mathcal{A}_{2})$. After taking these actions, they receive a shared immediate reward $r:\mathcal{S}\times\mathcal{A}_{1}\times\mathcal{A}_{2}\to\mathbb{R}$. The system then transitions to the next state $s_{t+1}$ according to the transition dynamics $\mathcal{T}(s_{t+1}\mid s_{t},a_{1,t},a_{2,t})\,\in\,\Delta(\mathcal{S})$ where $\Delta(\mathcal{S})$ denotes the set of probability distributions over $\mathcal{S}$. The parameter $\gamma\in[0,1)$ is the discount factor that weights immediate versus future rewards. Given two policies $\pi_{1}$ and $\pi_{2}$, mapping each state to a distribution over possible actions for Agents 1 and 2, respectively, the expected return is captured by the value function $V(\pi_{1},\pi_{2})\;=\;\mathbb{E}\Bigl{[}\sum_{t=0}^{\infty}\gamma^{t}\,r(s_{t},a_{1,t},a_{2,t})\Bigr{]}$ which sums the discounted rewards from each timestep. We will use $V(\pi_{1},\pi_{2})$ in later sections to discuss how agents can maximize their joint return.
@@
Generative model of cooperation strategies: Following Liang et al. [18], we train a Variational Auto-Encoder (VAE) [16] on diverse cooperative strategies. The VAE is trained to reconstruct coordination trajectories from a dataset $D$, which is obtained by using simulated PBT agents to interact with each other in the environment. Here, we use Maximum Entropy Population [43] to provide the simulated partner agents, since MEP was shown to provide the best performance in Liang et al. [18]. After producing the dataset $D$ using MEP agents, we then train the generative model, specifically the VAE, on this data. The VAE consists of an encoder $q(z|\tau;\phi)$, which maps a trajectory $\tau=(s_{0},a_{0},...,s_{T})$ to a latent space, where $\phi$ represents the parameters of the decoder. The decoder $p(a_{t}|z,\tau_{0..t-1};\phi)$ predicts actions $a_{t}$ autoregressively using the latent $z$ and the agent’s history $\tau_{0..t-1}$. The VAE model is optimized using the ELBO loss described in [16] as
$\mathcal{L}(\theta,\phi)=\mathbb{E}_{\tau\sim D}\Bigg{[}\mathbb{E}_{z\sim q(\cdot|\tau,\phi)}\left[\sum_{t=1}^{T}\log p(a_{t}|z,\tau_{0..t-1};\phi)\right]-\beta\cdot KL\big{(}q(z|\tau,\phi)\,\|\,\mathcal{N}(0,I)\big{)}\Bigg{]}$
The $\beta$ parameter balances reconstruction accuracy against the KL divergence in the ELBO loss, ensuring the latent space resembles a standard normal distribution $\mathcal{N}(0,I)$. This is important, because we can then sample new cooperation partners by sampling a $z\sim\mathcal{N}(0,I)$, and using the decoder to obtain a partner policy $\pi_{P}:=p(a_{t}|z,\tau_{0..t-1};\phi)$. The choice of $\beta$ leads to a tradeoff between reconstruction quality and encoding disentangled latent representations [1]. During training, the $\beta$ term was scheduled to automatically save different versions of KL-regularized VAE models. This method by Liang et al. [18] has proved efficient in capturing diverse strategies and enabling strategy generation that is valid and cooperative.

### v4

@@ Abstract
Being able to cooperate with diverse humans is an important component of many economically valuable AI tasks, from household robotics to autonomous driving. However, generalizing to novel humans requires training on data that captures the diversity of human behaviors. Adversarial training is a promising method that allows dynamic data generation and ensures that agents are robust. It creates a feedback loop where the agent’s performance influences the generation of new adversarial data, which can be used immediately to train the agent. However, adversarial training is difficult to apply in a cooperative task; how can we train an adversarial cooperator? We propose a novel strategy that combines a pre-trained generative model to simulate valid cooperative agent policies with adversarial training to maximize regret. We call our method GOAT: Generative Online Adversarial Training. In this framework, GOAT dynamically searches the latent space of the generative model for coordination strategies where the learning policy—the Cooperator agent—underperforms. GOAT enables better generalization by exposing the Cooperator to various challenging interaction scenarios. We maintain realistic coordination strategies by keeping the generative model frozen, thus avoiding adversarial exploitation. We evaluate GOAT with real human partners, and the results demonstrate state-of-the-art performance on the Overcooked benchmark, highlighting its effectiveness in generalizing to diverse human behaviors.  [^1] Please check out our live interactive demo, game play videos, and training code at our website https://sites.google.com/view/goat-2025/home
@@ 1 Introduction
In multi-agent human-AI cooperation (Carroll et al., 2020), training a cooperative agent to generalize across diverse human behaviors remains a formidable challenge. Distribution shift, limited training data, and highly dynamic human decision-making make training these systems difficult. Without adequate coverage of all possible human interactions, AI agents learn brittle policies and struggle to generalize to novel, unseen behaviors in real human cooperation partners.
Figure 1: Adversarial training framework for cooperative agents. (Left) A generative model encodes simulated agents into a latent space to learn diverse agent strategies, which are then used to generate different types of training partners. (Right) GOAT samples new partners to maximize the Cooperator agent’s regret, defined as the performance gap between self-play (the partner playing with itself) and cross-play (the partner playing with the Cooperator). The key idea is that the adversarial objective is constrained by the frozen generative model, which prevents it from generating self-sabotaging partners. But by applying regret-based adversarial training to search over the policies that can be generated by the model, we can expose the Cooperator to a curriculum of challenging training partners, ensuring it is robust to interacting with diverse human partners at test time.
@@
Adversarial training (Tucker et al., 2020; Cui et al., 2023; Fujimoto & Pedersen, 2022; Rutherford et al., 2024) could provide a computationally efficient solution for training robust Cooperators by exposing both the Cooperator agent’s limitations and underexplored areas of the partner strategy space. It can be used to automatically generate a curriculum of challenging tasks targeted to the learner’s weaknesses, resulting in a more robust policy Dennis et al. (2021). Yet despite its potential, adversarial training within cooperative AI introduces unique challenges; what does it mean to train an adversarial cooperator? In a single-agent task, Adversary’s sole objective is to challenge the learning agent as much as possible by minimizing its performance. However, in a collaborative task, the adversarial agent has to find weaknesses in the learner while still simulating a realistic cooperative partner policy. If the adversary is not constrained to such policies, then it can become overly combative, resorting to strategies that sabotage the cooperation task, and is thus no longer a valid simulation of what a human cooperation partner might do. The challenge lies in simulating realistic yet challenging partners that maintain cooperative intentions while simultaneously correcting weaknesses in the AI agent’s policy.
For single-agent tasks, techniques like Unsupervised Environment Design (UED) (Dennis et al., 2021; Parker-Holder et al., 2022; Jiang et al., 2021) constrains the Adversary so that it can only generate valid tasks for which there is a solution. This is accomplished by maximizing the learning agent’s regret, which is the performance gap between the optimal agent and the learning agent. The regret objective proves effective because it challenges the learning agent with a curriculum of increasingly difficult, yet still feasible tasks. In the cooperative setting, attempts have been made to create a regret-like objective (Charakorn et al., 2023; Villin et al., 2025; Rutherford et al., 2024; Wang et al., 2025; Erlebach & Cook, 2024). In methods like Charakorn et al. (2023), a diverse population is trained by optimizing for the difference between self-play and cross-play performance. However, naively applying this type of adversarial objective can lead to policies learning to sabotage the game. While attempts have been made to add additional objectives to counterbalance this effect Sarkar et al. (2023), they have not been able to address the issue that minimizing cross-play inherently incentivizes overly adversarial, uncooperative behavior.
@@
How can we ensure that the adversarial agent can only generate cooperative behaviors that do not engage in self-sabotage but can still fully explore the space of challenging partner strategies? We propose a novel adversarial training method to achieve these criteria. We first pre-train a generative model of partner strategies on diverse cooperation data, following Liang et al. (2024). We then place this generative model into an adversarial training loop, where we train an Adversary to search for embedding vectors that, when passed through the generative model, decode into simulated partners that maximize the Cooperator’s regret. Because we only update the weights of the Adversary and not the generative model, we ensure that all generated partners are cooperative, thereby addressing the intentional sabotage seen in adversarial training. In GOAT, the Adversary is motivated to maximize the regret, searching for valid coordination strategies where the Cooperator agent underperforms. In contrast, the Cooperator attempts to adapt to the Adversary and minimize the regret. Unlike previous adversarial approaches that risk learning degenerate strategies, this cyclic training method provides a natural curriculum for the Cooperator agent and unlocks the benefits of adversarial training for robustly generalizing to cooperate with diverse partners.
@@
We evaluate GOAT on three distinct domains: 1) One-Step Cooperative Matrix Game (CMG); 2) Cooperative Reaching Game (CRG); and 3) Overcooked (Carroll et al., 2020), a popular human-AI zero-shot coordination benchmark. We compare to 5 competitive baselines (Zhao et al., 2022; Sarkar et al., 2023; Liang et al., 2024; Strouse et al., 2022; Carroll et al., 2020) and show that GOAT achieves improved performance across all three popular environments. In the Overcooked environment, we tested GOAT live against real, novel human partners and found that it significantly improves cooperation performance.
@@
  • We introduce a novel adversarial training method for zero-shot coordination that trains both the Adversary and Cooperator agents online in a common payoff game, resulting in a dynamic curriculum where the Adversary searches over the space of partner policies that a generative model can simulate. The generative model is limited to simulating cooperative partners, which ensures the adversarial training process generates challenging but realistic cooperative partners.
  • We outperform 5 competitive baselines across 3 distinct popular cooperative environments.
  • We perform a real-time evaluation with novel human partners, using the popular Overcooked benchmark (Carroll et al., 2020) commonly used in human-AI cooperation research like Strouse et al. (2022); Charakorn et al. (2023); Zhao et al. (2022); Sarkar et al. (2023); Liang et al. (2024). As compared to recent high-scoring techniques, we show that our method leads to the best cooperation performance with diverse people.
@@ 2 Related Work
Zero-shot coordination (ZSC). Human-AI coordination requires training cooperative agents that can generalize to interactions with different humans in zero-shot settings, and thus navigate a vast coordination behavior space. (Hu et al., 2020; Kirk et al., 2023). Behavior cloning (Carroll et al., 2020) is limited by expensive and often scarce human data. Self-Play (SP) (Samuel, 1959; Silver et al., 2017; 2018; OpenAI et al., 2019) is an automatic training algorithm that does not need human data but converges to rigid policies that generalize poorly to novel human partners. Population-Based Training (PBT) (Jaderberg et al., 2017; Vinyals et al., 2019; Parker-Holder et al., 2020; Jung et al., 2020) has become a popular automatic training paradigm for ZSC, in which a Cooperator policy is trained using a large population of simulated partners (Hong et al., 2018; Strouse et al., 2022; Lupu et al., 2021; Charakorn et al., 2023; Zhao et al., 2022; Sarkar et al., 2023).
@@
Population diversity objectives. In an effort to ensure the simulated agent population covers diverse human strategies, various methods (Hong et al., 2018; Parker-Holder et al., 2020; Wu et al., 2023; Lupu et al., 2021; Zhao et al., 2022; Charakorn et al., 2023; Sarkar et al., 2023; Rahman et al., 2023; Yuan et al., 2023) improve PBT by optimizing the diversity of the agents in the population. The population can be created by reward-shaping (Tang et al., 2021; Yu et al., 2023), manual design (Ghosh et al., 2020; Wang et al., 2022; Xie et al., 2021), and quality diversity (Canaan et al., 2019; Pugh et al., 2016; Wu et al., 2023; Fontaine & Nikolaidis, 2021). However, these methods typically use domain knowledge and fail to generalize and scale. Statistical methods like (Hong et al., 2018; Eysenbach et al., 2018; Parker-Holder et al., 2020; Derek & Isola, 2021; Lupu et al., 2021; Zhao et al., 2022) automate the diverse population generation, but Sarkar et al. (2023) demonstrated that trajectory distribution methods can result in minor variations in the policies rather than behaviorally different policies.
@@
Min-Max optimization. Another line of work maximizes diversity by training self-play policies while minimizing cross-play rewards with the policies in the population (Cui et al., 2023; Charakorn et al., 2023; Rahman et al., 2023; 2024; Villin et al., 2025; Sarkar et al., 2023). However, agents might learn to sabotage the cross-play game while retaining high self-play performance with handshaking protocols, such that if their partner fails to complete the handshake correctly, they throw the game. To address this, Sarkar et al. (2023) introduces mixed-play, balancing self-play and cross-play by randomly sampling initial game states from both. However, such a solution still leads to rigid policies that fail to generalize, and does not address the fundamental problem that training a partner to minimize cross-play incentives sabotage. In this work, we only sample partners from a frozen, pre-trained generative model; since we do not train the partner policy on the adversarial objective, we prevent handshaking and sabotage. In addition, unlike prior work on this topic, we use online adversarial training to find and exploit weaknesses in the cooperator, instead of using adversarial objectives to pre-train a diverse partner population.
@@
Generative models for Cooperation. Derek & Isola (2021); Wang et al. (2023); Liang et al. (2024) show that generative models can be trained on policies to simulate diverse and valid collaborative policies. Recently, Liang et al. (2024) proposed training a Variational Autoencoder (Kingma & Welling, 2022), on simulated cooperative agent trajectories to generate synthetic training data to improve human-AI coordination. This approach outperforms previous methods because the generative model enables sampling random novel partner strategies that interpolate between or combine strategies from the data. GOAT leverages this approach to proactively generate challenging novel partners strategies to train the cooperator robustly.
@@ 3 Preliminaries
Two-player Markov game: We consider the two-player Markov game (Littman, 1994) as a tuple $\bigl(\,\mathcal{S},\,\mathcal{A}_{1},\,\mathcal{A}_{2},\,r,\;\mathcal{T},\;\gamma\bigr)$ where $\mathcal{S}$ is the state space shared by all the agents. $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ denote the action space of agents such that at each time step in the environment, and each agent selects one action $a_{1,t}\in\mathcal{A}_{1}$, $a_{2,t}\in\mathcal{A}_{2}$ sampled from their own action space according to their policies $\pi_{1}:\mathcal{S}\to\Delta(\mathcal{A}_{1})$ and $\pi_{2}:\mathcal{S}\to\Delta(\mathcal{A}_{2})$. After taking these actions, they receive a shared immediate reward $r:\mathcal{S}\times\mathcal{A}_{1}\times\mathcal{A}_{2}\to\mathbb{R}$. The system then transitions to the next state $s_{t+1}$ according to the transition dynamics $\mathcal{T}(s_{t+1}\mid s_{t},a_{1,t},a_{2,t})\,\in\,\Delta(\mathcal{S})$ where $\Delta(\mathcal{S})$ denotes the set of probability distributions over $\mathcal{S}$. The parameter $\gamma\in[0,1)$ is the discount factor that weights immediate versus future rewards. The expected joint return is captured by the value function $V(\pi_{1},\pi_{2})\;=\;\mathbb{E}\Bigl[\sum_{t=0}^{\infty}\gamma^{t}\,r(s_{t},a_{1,t},a_{2,t})\Bigr]$ which sums the discounted rewards from each timestep.
@@
Generative model of cooperation strategies: Following Liang et al. (2024), we train a Variational Auto-Encoder (VAE) (Kingma & Welling, 2022) on diverse cooperative strategies. The VAE is trained to reconstruct the partner’s actions from a dataset of coordination trajectories $D$, which is obtained from interactions between simulated PBT (Population-Based Training) agents within the environment. Here, we use a fixed agent population with training methods like Maximum Entropy Population (Zhao et al., 2022) and CoMeDi (Sarkar et al., 2023) to provide the simulated partner agents. The VAE consists of an encoder $q(z|\tau;\phi)$, which maps a trajectory $\tau=(s_{0},a_{0},...,s_{T})$ to a latent space, where $\phi$ represents the parameters of the decoder. The decoder $p(a_{t}|z,\tau_{0..t-1};\phi)$ predicts actions $a_{t}$ autoregressively using the latent $z$ and the agent’s history $\tau_{0..t-1}$. The VAE model is optimized using the ELBO loss described in (Kingma & Welling, 2022). The $\beta$ parameter that balances reconstruction accuracy against the KL divergence in the ELBO loss, ensuring the latent space resembles a standard normal distribution $\mathcal{N}(0,I)$. This is important for our method, because it enables generating new cooperation partners by sampling a latent $z\sim\mathcal{N}(0,I)$, and using the decoder to obtain a partner policy $\pi_{P}:=p(a_{t}|z,\tau_{0..t-1};\phi)$. The choice of $\beta$ leads to a tradeoff between reconstruction quality and encoding disentangled latent representations (Burgess et al., 2018).